OpenStudy (anonymous):

Solve the system of equations: x + y = 5 x – y = -9

OpenStudy (angela210793):

2x=-4 x=-2 -2+y=5 y=7

OpenStudy (anonymous):

add the equations x+y+x-y=5-9 => 2x=-4 -> x=-2. putting value of x in equation 1 -2+y=5 y=7;